# Adam Exploits $\ell_{\infty}$-geometry of Loss Landscape via Coordinate-wise Adaptivity

### v1

@@ 4.2 GPT-2 on language modeling task
We train GPT-2 small (124M parameters) [^3] We use nanoGPT codebase, https://github.com/karpathy/nanoGPT on the OpenWebText corpus containing more than 9B tokens for 100k iterations with sequence length of $512$ sequence length and $480$ sentences per batch. We use cosine learning rate schedule of the same peak learning rate $6\times 10^{-4}$, which is also the default of nanoGPT codebase. We did a grid search to find the maximum possible peak learning rate for $\mathtt{SGD}$  [^4] We tried $0.00001,0.00003,0.0001,0.0003,0.001,0.003,0.01,0.03,0.1,0.3,1$. . The training losses and evaluation losses of different optimizers are plotted in Figure 1. As mentioned in Section 1, $\mathtt{Adam}$ converges faster than $\mathtt{AdaSGD}$ while they both converge faster than $\mathtt{rotated\ Adam}$. Since we propose the $(1,1)$-norm of Hessian as a non rotation-invariant metric that can affect the convergence rate of $\mathtt{Adam}$, we also measure it for original loss function $L$ and rotated loss function $\tilde{L}$ on checkpoints trained with different losses. The results are presented in Table 2. The same correlation between norms and convergence rates holds here. The smaller the norm is, the faster the optimizer works.
Optimizer Smoothness Metric Matrix Norm Estimated Value
AdaSGD $H(L,\Phi_{\mathtt{AdaSGD}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$ $24.86$
Adam $H(L,\Phi_{\mathtt{Adam}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}$ $3.2$
Rotated Adam $H(L\circ\mathcal{R},\Phi_{\mathtt{Adam}})$ $\left\|R^{\top}\nabla^{2}L({\bm{x}})R\right\|_{1,1}$ $36.16$
@@
Due to the huge parameter count, it is computationally infeasible to sample a random orthogonal matrix from Haar measure over the orthogonal group. Thus we decide to sample from an alternative distribution over orthogonal matrices, where the final rotation matrix is the composition of a sequence of simple orthogonal matrices, including random permutation and apply random orthogonal transformation on the both side of matrix-valued parameters. See details in Section D.1.
@@
We first subsample a fixed batch of training data for estimating the $(1,1)$-norm of Hessian matrix. The high-level idea here is to compute the Hessian vector product between Hessian of training loss on this batch and a sequence of random Cauchy vectors and take the $\ell_{1}$ norm of the coordinate-wise median of the resulting sequence of Hessian vector products. Because Cauchy distribution is 1-stable, the resulting product is also a vector of Cauchy random variables, and each of them has the magnitude equal to $\ell_{1}$ norm of the corresponding row of the Hessian. Thus when we have infinitely many samples, the $\ell_{1}$ norm of the coordinate-wise median converges almost surely to the $(1,1)$-norm of Hessian. See details in Section D.2, where we also provide a non-asymptotic high-probability multiplicative bound for the estimation error.
@@ 4.3 ResNet18 on CIFAR-10
To further test the correlation between $\Phi$-smoothness and the optimization performance holds for architectures other than transformers, we conduct a similar experiment on ResNet18 trained on CIFAR-10 (Krizhevsky & Hinton, 2009). We applied random crop and random horizontal flip augmentations over the training data to promote better generalization. We tuned each optimizer through searching over the same grid of learning rates [^5] We used the following values: $6.25\times 10^{-4}$, $1.25\times 10^{-3}$, $2.5\times 10^{-3}$, $5.0\times 10^{-3}$, $1.0\times 10^{-2}$, $2.0\times 10^{-2}$, $4.0\times 10^{-2}$, $8.0\times 10^{-2}$, $1.6\times 10^{-1}$, $3.2\times 10^{-1}$, $6.4\times 10^{-1}$, $1.28\times 10^{0}$. The number of iterations is adjusted per batch size to result in  20 epochs for each training run (for instance, 4000 iterations were used for a batch size of 256, and 1000 iterations were used for a batch size of 1024). For the training loss and training accuracy plotted in Figure 2, it is measured on a subset of augmented training data that is the same size of evaluation set. The evaluation loss and accuracy are measured on the entire evaluation set without the augmentation. Track running stats is set to false at initialization.
@@
Figure 2 depicts the loss and accuracy curves for the best performing hyperparameters chosen over the training set’s final loss for batch size $256$. [^6] We have intentionally limited the number of training iterations to emphasize the difference of optimizers in terms of training speed over generalization. We also provide the results for other choices of batch size in Table 3. When it comes to optimization speed, even for ResNet18, $\mathtt{Adam}$ is always better than and they are always better than $\mathtt{AdaSGD}$, $\mathtt{SGD}$, and $\mathtt{rotated\ Adam}$ across different batch sizes. Note that this does not contradict with common practice of training ResNet with $\mathtt{SGD}$, where the main goal is to get better generalization and the training budget is large so all optimizers can easily achieve full training accuracy. In our experiment, we study optimization speed and intentionally limit the number of steps.
@@
We also measure the Hessian for checkpoints obtained at batch size $256$ and the results are in Table 4. The correlation between norms and convergence rates still holds here. When the $(1,1)$-norm divided by $d$ is smaller than spectral norm, $\mathtt{Adam}$ optimizes faster than $\mathtt{AdaSGD}$.
Optimizer Smoothness Metric Matrix Norm Estimated Value
AdaSGD $H(L,\Phi_{\mathtt{AdaSGD}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$ $1.5355$
@@ 5 Related Works
Previous work tries to analyze the difference between $\mathtt{Adam}$ and $\mathtt{SGD}$ from different perspectives. Zhou et al. (2018) proves a faster convergence rate of $\mathtt{Adam}$ than $\mathtt{SGD}$ when the stochastic gradients are sparse. Zhang et al. (2020) suggests that $\mathtt{SGD}$ suffers more from heavy-tailed noise than $\mathtt{Adam}$. Pan & Li (2023) claims that $\mathtt{Adam}$ has lower directional sharpness because of the effect of coordinate-wise clipping. Other works also consider the coordinate-wise normalization of $\mathtt{Adam}$ (Balles & Hennig, 2018; Kunstner et al., 2022). Kunstner et al. (2024) shows that the heavy-tailed class imbalance in language modeling tasks will cause $\mathtt{SGD}$ to converge slower when it can only optimize majority class well. Zhang et al. (2024a) finds that $\mathtt{Adam}$ is better at handling the block heterogeneity of Hessian matrix, which is a specific phenomenon in transformers. When viewing $\mathtt{Adam}$ as an adaptive method, there are works showing that adaptive methods have an advantage of achieving optimal convergence rate without relying on problem-dependent constant (Ward et al., 2020; Levy et al., 2021).
@@
There are many works showing convergence rate for $\mathtt{Adam}$ (Zhou et al., 2018; Chen et al., 2018; Zou et al., 2019; Shi & Li, 2021; Guo et al., 2021; Défossez et al., 2022; Zhang et al., 2022b). Most of them rely on the smoothness of the loss function, which is measured w.r.t. $\ell_{2}$ norm. Zhang et al. (2019) proposes the $(L_{0},L_{1})$ smoothness condition should be more reasonable than globally bounded smoothness. Li et al. (2024) further generalizes the $(L_{0},L_{1})$ smoothness condition. However, they still focus on the default $\ell_{2}$ norm which is rotation-invariant. To the best of our knowledge, we are the first to assume gradient Lipschitzness under $\ell_{\infty}$ norm for the analysis on $\mathtt{Adam}$.
@@ 6 Conclusion
We give a new convergence analysis (Theorem 3.5) for $\mathtt{Adam}$ in the stochastic non-convex setting using a novel smoothness assumption. We show the convergence rate for the 1-norm of the gradient is $O\left(\frac{1}{\sqrt{T}}\right)$ in the deterministic case (Corollaries 3.7 and 3.7) and $O\left(\left(\frac{\log T}{T}\right)^{1/4}\right)$ in the stochastic case (Corollary 3.6). We also extend our analysis to blockwise $\mathtt{Adam}$ on loss $L$ with respect to an arbitrary partition of the parameters $\Phi$ (Theorem 3.11) using the corresponding smoothness $H(L,\Phi)$ (Definition 3.9). Our bound for $\mathtt{Adam}$ involves $(1,1)$-norm of Hessian, rather than the spectral norm of Hessian, which is relevant to the convergence speed of $\mathtt{AdaSGD}$. This leads to significantly better smoothness conditions for deep learning models including ResNet-18 and GPT2 empirically. Our experiments also verify that the smoothness measure $H(L,\Phi)$ positively correlates with the optimization speed of blockwise $\mathtt{Adam}$ with respect to the partition $\Phi$.
@@ Acknowledgement
The authors would like to thank Khashayar Gatmiry and Sharut Gupta for helpful discussion and preliminary experiments in exploring the idea of rotated Adam. ZL is supported by OpenAI.
Appendix A Convergence rate of $\mathtt{SignGD}$ for deterministic loss
@@
Appendix B Invariance property of $\mathtt{Adam}$ and $\mathtt{SGD}$
See 2.2
Proof of Theorem 2.2.
For $\mathtt{SGD}$ and $\mathtt{AdaSGD}$, we will show they are rotation-invariant by induction. For any rotating transformation $\mathcal{R}({\bm{x}})=\bm{R}{\bm{x}}$, suppose $\tilde{{\bm{x}}}_{s}=\mathcal{R}^{-1}({\bm{x}}_{s})=\bm{R}^{\top}{\bm{x}}_{s}$ holds for $s\leq t-1$. Then we have that $\tilde{{\bm{g}}}_{t}=\nabla_{\tilde{{\bm{x}}}}\tilde{L}_{t}(\tilde{{\bm{x}}}_{t})=\bm{R}^{\top}\nabla_{{\bm{x}}}L(\bm{R}^{-1}\tilde{{\bm{x}}}_{t-1})=\bm{R}^{\top}\nabla_{{\bm{x}}}L({\bm{x}}_{t-1})=\bm{R}^{\top}{\bm{g}}_{t}$ and $\tilde{{\bm{m}}}_{t}=\bm{R}^{\top}{\bm{m}}_{t}$. From the update rule of $\mathtt{SGD}$, we have that $\tilde{{\bm{x}}}_{t}=\tilde{{\bm{x}}}_{t-1}-\eta_{t}\tilde{{\bm{m}}}_{t}=\bm{R}^{\top}{\bm{x}}_{t-1}-\eta_{t}\bm{R}^{\top}{\bm{m}}_{t}=\bm{R}^{\top}({\bm{x}}_{t-1}-\eta_{t}{\bm{m}}_{t})=\bm{R}^{\top}{\bm{x}}_{t}$. For the update rule of $\mathtt{AdaSGD}$, we further have that $\left\|\tilde{{\bm{g}}}_{t}\right\|_{2}^{2}=\left\|{\bm{g}}_{t}\right\|_{2}^{2}$ because $\bm{R}$ is an orthogonal matrix. Then $\tilde{v}_{t}=v_{t}$ and the derivation is similar.
For $\mathtt{Adam}$ and $\mathtt{SignGD}$, it is easy to show by induction they are invariant w.r.t. any permutating transformation because the operation on gradient is performed on each coordinate separately. We only need to show they are not invariant w.r.t. a rotating transformation. We choose $\bm{R}=[\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}};\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}]$, $L_{t}({\bm{x}})=L({\bm{x}})=2x_{1}^{2}+x_{2}^{2}$. Due to the update rule of $\mathtt{SignGD}$, it can only update ${\bm{x}}$ and $\tilde{{\bm{x}}}$ in the direction of $[1,1]$ and $[1,-1]$. But when rotating the update direction on $\tilde{{\bm{x}}}$ back to the space of ${\bm{x}}$. The update direction can only be $[1,0]$ or $[0,1]$ that are different from the update direction in the original space. Because the first step in $\mathtt{Adam}$ takes the same direction in $\mathtt{SignGD}$, we simultaneously show that both $\mathtt{SignGD}$ and $\mathtt{Adam}$ are not rotation-invariant. ∎
Appendix C Proof Details
C.1 Proof for convergence rate of blockwise $\mathtt{Adam}$
We also need new notations and different assumptions for specific partition $\Phi(\cdot)$. For specific block $b$, $(b)$ is defined as $\Phi^{-1}(b)=\{i|\Phi(i)=b\}$ and $d_{b}=\#(b)$, the number of parameters in block $b$.
As mentioned in Section 3.4, we will use Lemma 3.12 to bound the second order term under smoothness Definition 3.9 and use Lemma 3.13 to better control the growth of the sum of second order term. See 3.12

### v3

@@ 4.2 GPT-2 on language modeling task
We train GPT-2 small (124M parameters) [^3] Our codebase is built upon nanoGPT codebase https://github.com/karpathy/nanoGPT. on the OpenWebText corpus containing more than 9B tokens for 100k iterations with sequence length of $512$ sequence length and $480$ sentences per batch. We use cosine learning rate schedule of the same peak learning rate $6\times 10^{-4}$ for all the adaptive optimizers, which is also the default of nanoGPT codebase. We did a grid search to find the maximum possible peak learning rate for $\mathtt{SGD}$  [^4] We tried $0.00001,0.00003,0.0001,0.0003,0.001,0.003,0.01,0.03,0.1,0.3,1$. . The training losses and evaluation losses of different optimizers are plotted in Figure 1. As mentioned in Section 1, $\mathtt{Adam}$ converges faster than $\mathtt{AdaSGD}$ while they both converge faster than $\mathtt{rotated\ Adam}$. Since we propose the $(1,1)$-norm of Hessian as a non-rotation-invariant metric that can affect the convergence rate of $\mathtt{Adam}$, we also measure it for the original loss function $L$ and rotated loss function $\tilde{L}$ on checkpoints trained with different losses. The results are presented in Table 2. The same correlation between norms and convergence rates holds here. The smaller the norm is, the faster the optimizer works.
Optimizer Smoothness Metric Upper Bound Estimated Value
AdaSGD $H(L,\Phi_{\mathtt{AdaSGD}})$ $d\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$ $4.2446$
Adam $H(L,\Phi_{\mathtt{Adam}})$ $\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}$ $2.3538$
Rotated Adam $H(L\circ\mathcal{R},\Phi_{\mathtt{Adam}})$ $\left\|R^{\top}\nabla^{2}L({\bm{x}})R\right\|_{1,1}$ $14.3745$
@@
We also explore how learning rate can affect the performance of different optimizers by using peak learning rate $3\times 10^{-4}$ and $1.8\times 10^{-3}$ for all the adaptive optimizers. The training losses are plotted in Figure 3. All the optimizers perform worse with smaller learning rate, which aligns with the common understanding that optimizers tend to work better with larger learning rate as long as the training is still stable. When a larger learning rate is used, the performance of $\mathtt{Adam}$ is improved but the performance of rotated $\mathtt{Adam}$ becomes worse than with the default learning rate. The training with $\mathtt{AdaSGD}$ even completely failed. This suggests that another advantage of $\mathtt{Adam}$ over $\mathtt{AdaSGD}$: it can maintain stable training at a larger learning rate, which is often beneficial to faster and more efficient convergence.
Figure 3: Training losses of $\mathtt{Adam}$, $\mathtt{AdaSGD}$ and rotated $\mathtt{Adam}$ on GPT-2 under different learning rates. All the optimizers perform worse with smaller learning rate $3\times 10^{-4}$. Only $\mathtt{Adam}$ will perform better with larger learning rate $1.8\times 10^{-3}$.
@@
Due to the huge parameter count, it is computationally infeasible to sample a random orthogonal matrix from Haar measure over the orthogonal group. Thus we decide to sample from an alternative distribution over orthogonal matrices, where the final rotation matrix is the composition of a sequence of simple orthogonal matrices, including random permutation and apply random orthogonal transformation on both sides of matrix-valued parameters. See details in Section D.1. Concurrent work Maes et al. (2024) also run $\mathtt{Adam}$ on a rotated loss. They use a different way to achieve global rotation efficiently and conduct extensive module-wise rotation experiments for a more fine-grained analysis.
@@
The algorithm is defined in Algorithm 4. We first subsample a fixed batch of training data for estimating the $(1,1)$-norm of Hessian matrix. The high-level idea is to compute the matrix vector products between Hessian of training loss on this batch and a sequence of random Cauchy vectors. Then we take the $\ell_{1}$ norm of the coordinate-wise median of the resulting sequence of Hessian vector products. Because the Cauchy distribution is 1-stable, the resulting product is also a vector of Cauchy random variables, and the magnitude of each element equals to $\ell_{1}$ norm of the corresponding row of the Hessian. Thus with infinitely many samples, the $\ell_{1}$ norm of the coordinate-wise median converges almost surely to the $(1,1)$-norm of the Hessian. Below we provide a non-asymptotic high-probability multiplicative bound for the estimation error which depends mildly on the dimension $d$. More explanations and the proof of Theorem 4.1 are in Section D.2. Concurrent work Maes et al. (2024) also proposed another algorithm to measure $(1,1)$-norm.
Theorem 4.1.
For the estimate in Algorithm 4 with $n$ Cauchy vectors, it holds that
$P\left(\left|\sum_{j=1}^{d}\textup{median}(\left|\mathbf{H}_{j,:}\right|)-\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}\right|\geq\epsilon\left\|\nabla^{2}L({\bm{x}})\right\|_{1,1}\right)\leq 2d\exp(-\frac{n\Delta^{2}}{2})+2\exp{\left(-\frac{2n\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}{\pi^{2}}\right)}$
for every $\epsilon,\Delta\in(0,1)$ when $n\geq\frac{\pi^{3}}{2\epsilon^{2}\cos^{4}((1+\Delta)\frac{\pi}{4})}$.
In other words, for any $\epsilon,\delta\in(0,1)$, we can use $n=\Omega(\ln d+\frac{1}{\epsilon^{2}}\ln\frac{1}{\delta})$ hessian-vector product of the loss $L$ at parameter ${\bm{x}}$ and $nd$ extra computation time to get an estimation of (1,1)-norm of $L$ with at most $\epsilon$ multiplicative error and at least probability $1-\delta$.
@@ 4.3 ResNet18 on CIFAR-10
To further test whether the correlation between $\Phi$-smoothness and the optimization performance holds for architectures other than transformers, we conduct a similar experiment on ResNet18 trained on CIFAR-10 (Krizhevsky & Hinton, 2009). We applied random crop and random horizontal flip augmentations over the training data to promote better generalization. We tuned each optimizer through searching over the same grid of learning rates [^5] We used the following values: $6.25\times 10^{-4}$, $1.25\times 10^{-3}$, $2.5\times 10^{-3}$, $5.0\times 10^{-3}$, $1.0\times 10^{-2}$, $2.0\times 10^{-2}$, $4.0\times 10^{-2}$, $8.0\times 10^{-2}$, $1.6\times 10^{-1}$, $3.2\times 10^{-1}$, $6.4\times 10^{-1}$, $1.28\times 10^{0}$. The number of iterations is adjusted per batch size to result in  20 epochs for each training run (for instance, 4000 iterations were used for a batch size of 256, and 1000 iterations were used for a batch size of 1024). For the training loss and training accuracy plotted in Figure 4, it is measured on a subset of augmented training data that is the same size of evaluation set. The evaluation loss and accuracy are measured on the entire evaluation set without the augmentation. Track running stats is set to false at initialization.
Figure 4 depicts the loss and accuracy curves for the best performing hyperparameters chosen over the training set’s final loss for batch size $256$. [^6] We have intentionally limited the number of training iterations to emphasize the difference of optimizers in terms of training speed over generalization. We also provide the results for other choices of batch size in Table 3. When it comes to optimization speed, even for ResNet18, $\mathtt{Adam}$ is always better than $\mathtt{rotated\ Adam}$ and they are always better than $\mathtt{AdaSGD}$ and $\mathtt{SGD}$ across different batch sizes. Note that this does not contradict with common practice of training ResNet with $\mathtt{SGD}$, where the main goal is to get better generalization and the training budget is large so all optimizers can easily achieve full training accuracy. In our experiment, we study optimization speed and intentionally limit the number of steps.
@@
We also measure the Hessian for checkpoints obtained at batch size $256$ and the results are in Table 4. The correlation between norms and convergence rates still holds here. When the $(1,1)$-norm is smaller than $d$ times spectral norm, $\mathtt{Adam}$ optimizes faster than $\mathtt{AdaSGD}$.
Optimizer Smoothness Metric Upper Bound Estimated Value
AdaSGD $H(L,\Phi_{\mathtt{AdaSGD}})$ $d\left\|\nabla^{2}L({\bm{x}})\right\|_{2}$ $1.5355$
@@ 5 Related Works
Previous work tries to analyze the difference between $\mathtt{Adam}$ and $\mathtt{SGD}$ from different perspectives. Zhou et al. (2018) proves a faster convergence rate of $\mathtt{Adam}$ than $\mathtt{SGD}$ when the stochastic gradients are sparse. Zhang et al. (2020) suggests that $\mathtt{SGD}$ suffers more from heavy-tailed noise than $\mathtt{Adam}$. Pan & Li (2023) claims that $\mathtt{Adam}$ has lower directional sharpness because of the effect of coordinate-wise clipping. Other works also consider the coordinate-wise normalization of $\mathtt{Adam}$ (Balles & Hennig, 2018; Kunstner et al., 2022). Kunstner et al. (2024) shows that the heavy-tailed class imbalance in language modeling tasks will cause $\mathtt{SGD}$ to converge slower when it can only optimize majority class well. Zhang et al. (2024a) finds that $\mathtt{Adam}$ is better at handling the block heterogeneity of Hessian matrix, which is a specific phenomenon in transformers. When viewing $\mathtt{Adam}$ as an adaptive method, there are works showing that adaptive methods have an advantage of achieving optimal convergence rate without relying on problem-dependent constant (Ward et al., 2020; Levy et al., 2021). Ling et al. (2022) focuses on the rotation equivariance property in the geometry optimization setting.
@@
There are many works showing convergence rate for $\mathtt{Adam}$ (Zhou et al., 2018; Chen et al., 2018; Zou et al., 2019; Shi & Li, 2021; Guo et al., 2021; Défossez et al., 2022; Zhang et al., 2022b). Most of them rely on the smoothness of the loss function, which is measured w.r.t. $\ell_{2}$ norm. Zhang et al. (2019) proposes the $(L_{0},L_{1})$ smoothness condition should be more reasonable than globally bounded smoothness. Li et al. (2024) further generalizes the $(L_{0},L_{1})$ smoothness condition. However, they still focus on the default $\ell_{2}$ norm which is rotation-invariant. Concurrent work Maes et al. (2024) claims analyzing $\mathtt{Adam}$ requires rotation dependent assumptions by conducting various rotating experiments. To the best of our knowledge, we are the first to assume gradient Lipschitzness under $\ell_{\infty}$ norm for the analysis on $\mathtt{Adam}$.
@@ 6 Conclusion
We give a new convergence analysis (Theorem 3.5) for $\mathtt{Adam}$ in the stochastic non-convex setting using a novel smoothness assumption. We show the convergence rate for the $\ell_{1}$ norm of the gradient is $O\left(\frac{1}{\sqrt{T}}\right)$ in the deterministic case (Corollaries 3.7 and 3.7) and $O\left(\left(\frac{\log T}{T}\right)^{1/4}\right)$ in the stochastic case (Corollary 3.6). We also extend our analysis to blockwise $\mathtt{Adam}$ on loss $L$ with respect to an arbitrary partition of the parameters $\Phi$ (Theorem 3.11) using the corresponding smoothness $H(L,\Phi)$ (Definition 3.12). Our bound for $\mathtt{Adam}$ involves $(1,1)$-norm of Hessian, rather than the spectral norm of Hessian, which is relevant to the convergence speed of $\mathtt{AdaSGD}$. This leads to significantly better smoothness conditions for deep learning models including ResNet-18 and GPT2 empirically. Our experiments also verify that the smoothness measure $H(L,\Phi)$ positively correlates with the optimization speed of blockwise $\mathtt{Adam}$ with respect to the partition $\Phi$.
@@ Acknowledgement
The authors would like to thank Khashayar Gatmiry and Sharut Gupta for the helpful discussion and preliminary experiments in exploring the idea of rotated Adam and Xiaoyu Chen for discussions on extending the analysis to blockwise Adam and steepest descent w.r.t. general $\Phi$-norm. ZL is supported by OpenAI Superalignment Grant.
Appendix A Equivariance Property of $\mathtt{Adam}$ and $\mathtt{SGD}$
See 2.2
Proof of Theorem 2.2.
For $\mathtt{SGD}$ and $\mathtt{AdaSGD}$, we will show they are rotation-equivariant by induction. For any rotating transformation $\mathcal{R}({\bm{x}})=\bm{R}{\bm{x}}$, suppose $\tilde{{\bm{x}}}_{s}=\mathcal{R}^{-1}({\bm{x}}_{s})=\bm{R}^{\top}{\bm{x}}_{s}$ holds for $s\leq t-1$. Then we have that $\tilde{{\bm{g}}}_{t}=\nabla_{\tilde{{\bm{x}}}}\tilde{L}_{t}(\tilde{{\bm{x}}}_{t})=\bm{R}^{\top}\nabla_{{\bm{x}}}L(\bm{R}^{-1}\tilde{{\bm{x}}}_{t-1})=\bm{R}^{\top}\nabla_{{\bm{x}}}L({\bm{x}}_{t-1})=\bm{R}^{\top}{\bm{g}}_{t}$ and $\tilde{{\bm{m}}}_{t}=\bm{R}^{\top}{\bm{m}}_{t}$. From the update rule of $\mathtt{SGD}$, we have that $\tilde{{\bm{x}}}_{t}=\tilde{{\bm{x}}}_{t-1}-\eta_{t}\tilde{{\bm{m}}}_{t}=\bm{R}^{\top}{\bm{x}}_{t-1}-\eta_{t}\bm{R}^{\top}{\bm{m}}_{t}=\bm{R}^{\top}({\bm{x}}_{t-1}-\eta_{t}{\bm{m}}_{t})=\bm{R}^{\top}{\bm{x}}_{t}$. For the update rule of $\mathtt{AdaSGD}$, we further have that $\left\|\tilde{{\bm{g}}}_{t}\right\|_{2}^{2}=\left\|{\bm{g}}_{t}\right\|_{2}^{2}$ because $\bm{R}$ is an orthogonal matrix. Then $\tilde{v}_{t}=v_{t}$ and the derivation is similar.
For $\mathtt{Adam}$ and $\mathtt{SignGD}$, it is easy to show by induction they are equivariant w.r.t. any permutating transformation because the operation on gradient is performed on each coordinate separately. We only need to show they are not equivariant w.r.t. a rotating transformation. We choose $\bm{R}=[\frac{1}{\sqrt{2}},\frac{1}{\sqrt{2}};\frac{1}{\sqrt{2}},-\frac{1}{\sqrt{2}}]$, $L_{t}({\bm{x}})=L({\bm{x}})=2x_{1}^{2}+x_{2}^{2}$. Due to the update rule of $\mathtt{SignGD}$, it can only update ${\bm{x}}$ and $\tilde{{\bm{x}}}$ in the direction of $[1,1]$ and $[1,-1]$. But when rotating the update direction on $\tilde{{\bm{x}}}$ back to the space of ${\bm{x}}$. The update direction can only be $[1,0]$ or $[0,1]$ that are different from the update direction in the original space. Because the first step in $\mathtt{Adam}$ takes the same direction in $\mathtt{SignGD}$, we simultaneously show that both $\mathtt{SignGD}$ and $\mathtt{Adam}$ are not rotation-equivariant. ∎
Appendix B Convergence rate of $\mathtt{SignGD}$ for deterministic loss
See 3.2
@@ Appendix B Convergence rate of $\mathtt{SignGD}$ for deterministic loss
Appendix C Proof for Convergence Rate of Blockwise $\mathtt{Adam}$
As mentioned in Section 3.3, we will use Lemma 3.13 to show the relationship between blockwisely-smoothness w.r.t. $\Phi$-norm and $H(L,\Phi)$. See 3.13